OpenStudy (anonymous):

The midpoint of A (-4, 2) and B(8, 5) is (2, 7/2) (-6, 3/2) (4, 7/2)

OpenStudy (kamibug):

Here's the formula for midpoint. :)

OpenStudy (igreen):

^^^ \(\dfrac{x_1+x_2}{2} + \dfrac{y_1+ y_2}{2}\) Plug in what we know: \(\dfrac{-4+8}{2} + \dfrac{2+ 5}{2}\) I think you can finish that. @ShaquanaBBurger

OpenStudy (anonymous):

thanks